OpenStudy (anonymous):

I will give a medal for the one who helps me answer my math problem the best please help me. Write the equation of the line that passes through the point (3,5) and is parallel to the line 2x+y=3. How do I know or figure out the line it is asking for?

OpenStudy (anonymous):

Firstly, you will want to simplify the line 2x+y=3 into an easy understandable form like y=mx+b

OpenStudy (anonymous):

okay so it will be y=-2x+3

OpenStudy (anonymous):

Once, you find that out, you should have the slope of that line. If you are trying to find a line parallel to the one given, you should know that two parallel lines will never cross each other so have the same slope

OpenStudy (anonymous):

Okay, so you have y=-2x+3, the new line will have the same slope as it is parallel and will never intersect with that line Next, you are given a point the new line passes (3,5) or (x,y). Your new equation should look like this so far y=-2x+b to find b plug in 3 for x and 5 for y and solve for b

OpenStudy (anonymous):

I got 5=-3 ?

OpenStudy (anonymous):

but the answer on the back is y=-2x+11 ... how?

OpenStudy (anonymous):

@Young-Lau

OpenStudy (anonymous):

so, when you plug in the points you get 5=-2(3)+b multiply -2*3 5=-6+b add 6 to both sides 11=b now you have b. Now put it into your equation to get y=-2x+11

OpenStudy (anonymous):

ohhhh !!!!!! I get it now :) thank you so much :) @Young-Lau

OpenStudy (anonymous):

You're welcome :)